### v4

@@ IV-C SPICE netlist verification
Our case study revealed that some generated netlists (Fig. 4 (➍)) exhibit issues such as floating nets. Although these issues may pass a simulator check, they pose significant risks for large-scale datasets where manual inspection is infeasible. To mitigate this, we developed a Python-based feedback mechanism (Fig. 4 (➎)) that enables LLMs to identify and self-correct such errors automatically, eliminating the need for manual intervention. This solution establishes a robust, closed-loop verification process, which can be further enhanced with additional netlist verification features. We are also releasing this as part of our open-source contribution [16].
@@ V-A Dataset Creation
Figure 8: Distribution of (a) Number of Components, (b) Number of Nodes, (c) Number of MOSFETs, and (d) Number of lines of SPICE code, respectively in the dataset.
@@
We curated a comprehensive dataset by selecting schematics from textbooks [24, 30]. Textbooks were chosen over research papers because they offer high-quality schematic images, clear context, and detailed explanations. This makes them more suitable for our purposes, as research papers often present high-level descriptions with advanced schematics that use abstract blocks for complex components. We collected over 2,100 schematic images and applied our proposed Masala-CHAI flow (Fig. 4) to generate the corresponding SPICE netlists. Alongside this, we collected captions and descriptions as detailed in Section IV. Fig. 7 demonstrates how Masala-CHAI translates a schematic into a SPICE netlist. Since our primary goal is to generate structurally correct netlists, each component is marked with default parameters.
@@
Our dataset is carefully characterized to showcase its diversity and complexity, as shown in Fig. 8: (a) displays the variation in the number of components within the schematics, reflecting the complexity of the circuits - a higher number indicates a more complex design; (b) illustrates the distribution of nodes in the schematics, which corresponds to the connectivity between components. Notably, accurately identifying and mapping MOSFET terminals to the correct nets is a significant challenge for LLMs; (c) focuses on the number of MOSFETs present, indicating increased complexity for SPICE generation as the number rises; (d) presents the distribution of the number of lines in the generated SPICE netlists, further emphasizing the varying levels of complexity encompassed in our dataset. By creating this diverse and challenging dataset, we provide a robust benchmark for evaluating and enhancing the capabilities of LLMs in automated SPICE netlist generation.
V-B Evaluation
@@
To evaluate the performance of our proposed method, we leverage graph-matching concepts to compare the structural similarity between SPICE netlists. Unlike traditional approaches, we parse the netlists to extract components and their connections. We identify components by their types (e.g., Resistor, Capacitor, MOSFET) to ensure that our graph representations remain invariant to varying component labels used during netlist generation. This results in a constructed graph “G”, where nodes represent circuit components and edges represent their interconnections, as illustrated Fig. 9. We use Graph Edit Distance (GED) as a quantitative metric to compare the representations of two SPICE netlists. GED ($G_{1}$,$G_{2}$) calculates the sum of the costs associated with transforming $G_{1}$ into $G_{2}$ through a series of edit operations on nodes and edges, such as insertions, deletions, and substitutions. To enhance the interpretability of our results, we normalize the GED to yield a final metric ranging from 0 to 100. This normalization is achieved by calculating the maximum possible GED, obtained by summing the total number of nodes and edges in both graphs and scaling accordingly. This standardized metric allows for a consistent and meaningful comparison across different circuits. Thus, the similarity score between the two netlists is:
@@
V-C Verifying against AMSNet [13] netlists
@@
We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset, which includes $\sim$800 circuit diagrams. Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Fig. 4). When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier. This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams. Fig. 9 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋). The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions. These findings strengthen Masala-CHAI accuracy and scalability in generating accurate netlists for various circuit diagrams.
V-D Finetuning
Finetuning entails adapting a pre-trained model to a specific downstream task [4, 3]. This process entails further training the model on a smaller task-specific dataset, enabling it to adjust its parameters to capture new patterns and subtle nuances relevant to the task without training from scratch. The advantages of fine-tuning are: 1) It substantially reduces training time and computational resources by leveraging the foundational knowledge already embedded in the model; 2) It enhances performance on specialized tasks, particularly when data is scarce; and 3) It facilitates the transfer of learned representations across different domains, thereby improving efficiency and effectiveness in model deployment.
GPT: OpenAI [31] released GPT models up to the state-of-the-art GPT-4o model, which is expensive. Fine-tuning [32] reduces cost and latency by replacing the expensive GPT-4o, with a fine-tuned GPT-4o-mini model. The fine-tuned GPT-4o-mini can achieve quality results similar to those of GPT-4o.
Table I reports preliminary results of fine-tuning the GPT’s models on the Masala-CHAI datasets generated. We picked 20 common design problems frequently used as a standalone circuit or sub-part of a bigger design. We prompted the LLM to generate (n = 10) samples of SPICE netlist for each circuit description and manually validated their correctness (the score reported is the number of correct SPICE netlists). We subdivided the 20 benchmarks into 3 different categories: Easy (ID=1-7), Medium (ID=8-14), and Hard (ID=15-20).  Table I provides a comprehensive evaluation of the performance of GPT-3.5-turbo and GPT-4o-mini, both in their ‘base’ and ‘finetune’ states, on a diverse set of circuit design tasks. Finetuning demonstrates significant performance gains across nearly all circuits, particularly for complex designs. For instance, in the (ID=8), GPT-3.5-turbo improves from a score of 4 to 10. At the same time, GPT-4o-mini achieves a perfect score of 10 after fine-tuning, demonstrating the impact of domain-specific training. This pattern is evident across other tasks such as the (ID=5), where fine-tuning boosts GPT-3.5-turbo from 5 to 9 and GPT-4o-mini from 6 to 10.
The results reveal an important trend: simple circuits, such as the (ID=1) and (ID=3), are well-handled even by the base models. More advanced circuits like the (ID=14) and (ID=17) require fine-tuning to achieve meaningful results. Notably, finetuning enables GPT-4o-mini to score 8 on the (ID=10), doubling its base score of 4, illustrating the model’s scalability for complex tasks. GPT-4o-mini consistently outperforms GPT-3.5-turbo in handling advanced designs, achieving superior scores across several tasks.
Table I reveals significant challenges, particularly with specialized circuits like the (ID=19) and (ID=20), where both models exhibit minimal gains even after finetuning.  Fig. 10 showcases representative examples of both successful and failed netlists generated by LLMs, illustrating critical design challenges and the effectiveness of our approach. Notably, in Fig. 10: (1) demonstrates the improper use of an NMOS as a current source rather than forming the intended cascode configuration; (2) highlights a fundamental error where the gate of a diode-connected transistor, expected to connect to the drain, is incorrectly connected to the source; (3) the complexity of a two-stage design further amplifies the difficulties in understanding and replicating the intended topology. These examples highlight the inherent limitations of unoptimized LLMs. Through the ‘Finetune’ GPT-4o-mini model, we successfully resolved these issues, achieving accurate connections and eliminating floating nets. This demonstrates the value of fine-tuning in enabling reliable and context-aware netlist generation.
Figure 10: Examples of passed and failed design cases generated by the fine-tuned model.
TABLE I: Results on 20 designs show the success rate of SPICE netlist generation (out of 10 samples, manually validated). ‘Base’ refers to the baseline model, and ‘Finetune’ indicates its fine-tuned version.
ID Circuit Description GPT-3.5-turbo GPT-4o-mini
Base Finetune Base Finetune
1 Common-source amplifier 8 10 8 10
2 2-stage common source amplifier with resistive load 6 10 8 10
3 Common-drain amplifier 8 10 10 10
4 common-gate amplifier 7 10 10 10
5 Single-Stage RC Low-Pass Filter 5 9 6 10
6 Source Degenerated Amplifier 6 8 6 10
7 Current Mirror 5 9 5 9
8 Common-source amplifier using active load 4 10 7 10
9 Cascode amplifier using NMOS and resistive load 6 10 6 8
10 1-stage differential amplifier 2 7 4 8
11 Diode-connected Amplifier 0 3 3 9
12 Buffer design using MOSFET 0 4 2 8
13 2-input NAND gate 0 5 0 8
14 2-stage amplifier with miller compensation 0 1 0 7
15 SRAM cell with 6 transistors 0 1 0 6
16 2-stage op-amp with differential inputs and single-handled output 0 0 0 4
17 Fully Differential Amplifier with Common-Mode Feedback 0 0 0 3
18 Cross-coupled LC oscillator 0 0 0 3
19 Telescopic cascode operational amplifier 0 0 0 2
20 Bandgap Reference Amplifier 0 0 0 0
VI Conclusion and Future Work
In this paper, we introduced Masala-CHAI, a novel framework to create large-scale datasets for SPICE netlists from documents comprising analog circuit schematics, addressing critical challenges in the design workflow. Through schematic preprocessing and prompt tuning, we developed a fully automated pipeline that leverages LLMs for scalable netlist generation. To the best of our knowledge, Masala-CHAI is the first LLM-based framework of its kind, and our released dataset and flow [16] provide a valuable resource for the community for future advancements.
Future work will focus on enhancing the framework with automated parameter tuning to ensure functionally validated netlists. Additionally, we aim to extend automation to the layout generation stage, addressing its inherent challenges by curating comprehensive datasets and fine-tuning LLMs for this task. These advancements will further streamline the analog design process, bridging the gap from schematic to GDSII and unlocking new efficiencies in IC design workflows.

### v5

@@ IV-C SPICE netlist verification
Our case study revealed that some generated netlists (Figure 5 (➎)) still exhibit issues such as floating nets, posing risks large-scale datasets where manual inspection is infeasible. To mitigate this, we introduced a final verification step (Figure 5 (➏)) that calls an LLM on the extracted SPICE netlist to automatically identify and correct common errors, for instance, removing floating nets as mentioned above. These checks are provided in our open-source code base, and can be further extended to address other sources of error, as required.
@@ V-A Dataset Creation
Figure 9: Distribution of (a) # Components, (b) # Nodes, (c) # MOSFETs, and (d) #lines of SPICE code in the dataset.
@@
We curated a comprehensive dataset by automating schematic extraction and SPICE netlist generation from 10 textbooks [29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. These textbooks were selected for their high-quality schematic images, clear context, and detailed explanations. After collecting $\sim$7,500 schematic images, we applied the Masala-CHAI framework (Figure 5) to generate SPICE netlists. Alongside, we collected captions and descriptions as detailed in Section IVFigure 8 demonstrates how Masala-CHAI translates a schematic into SPICE. Since our goal is to generate structurally correct netlists, each component is marked with default parameters.
@@
Our dataset is characterized to highlight its diversity and complexity in  Figure 9: (a) shows the variation in the number of components within schematics, where a higher count indicates greater circuit complexity; (b) presents the distribution of nodes, representing component connectivity. Accurately identifying and mapping MOSFET terminals to the correct nets is a significant challenge for LLMs; (c) focuses on the number of MOSFETs, where an increase in count corresponds to higher complexity in SPICE netlist generation; Finally, (d) illustrates the distribution of SPICE netlist line counts, emphasizing the circuit complexities. This characterization provides insights into the dataset, aiding in the evaluation of automated SPICE generation techniques. By curating this diverse and complex dataset, we establish a robust benchmark for evaluating and improving LLM performance in automated SPICE netlist generation. For more analysis see Appendix A-A and  A-B.
TABLE II: Evaluation benchmarks. We refer each design by its ID.
ID Circuit ID Circuit ID Circuit
1 Common-source amplifier 2 2-stage common src amplifier w/ res. load 3 Common-drain amplifier
4 Common-gate amplifier 5 Single-Stage RC Low-Pass Filter 6 Src Degenerated Amplifier
7 Current Mirror 8 Common-source amplifier using active load 9 Cascode amplifier using NMOS and res. load
10 1-stage differential amplifier 11 Diode-connected Amplifier 12 Buffer design using MOSFET
13 2-input NAND gate 14 2-stage op-amp with Miller compensation 15 SRAM cell with 6 transistors
16 2-stage op-amp with diff. i/p and single-handled o/p 17 Fully Diff. Amplifier w/ Common-Mode Feedback 18 Cross-coupled LC oscillator
19 Telescopic cascode op amplifier 20 Bandgap Reference Amplifier
TABLE III: Results on LLMs and Finetuned (‘FT’) variants on MASALA-CHAI dataset. ‘CL’ is CodeLlama. ‘DS’ is DeepSeek.
ID CL-70B CL-70B+FT DS-34B DS-34B + FT GPT-3.5 GPT-3.5 + FT GPT-4o GPT-4o + FT
Pass @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5 @1 @5
1 13.3 57.1 26.7 84.6 26.7 84.6 40 95.8 66.7 100 93.3 100 66.7 100 100 100
2 0.0 0.0 13.3 57.1 6.7 33.3 26.7 84.6 53.3 99.3 86.7 100 66.7 100 100 100
3 26.7 84.6 33.3 91.6 13.3 57.1 46.7 98.1 66.7 100 100 100 80 100 100 100
4 13.3 57.1 33.3 91.6 26.7 84.6 33.3 91.6 60 99.8 80 100 60 99.8 100 100
5 6.7 33.3 26.7 84.6 6.7 33.3 26.7 84.6 46.7 98.1 86.7 100 60 99.8 100 100
6 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 60.0 99.8 73.3 100 80 100 100 100
7 0.0 0.0 6.7 33.3 0.0 0.0 26.7 84.6 46.7 98.1 86.7 100 53.3 99.3 93.3 100
8 0.0 0.0 0.0 0.0 0.0 0.0 6.7 33.3 40.0 95.8 66.7 100 46.7 98.1 100 100
9 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 46.7 98.1 73.3 100 53.3 99.3 86.7 100
10 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 26.7 84.6 86.7 100 46.7 98.1 93.3 100
11 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 66.7 100 40 95.8 100 100
12 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 66.7 100 33.3 91.6 86.7 100
13 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 53.3 99.3 20.0 73.6 80 100
14 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 40 95.8 0.0 0.0 73.3 100
15 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 33.3 91.6 0.0 0.0 86.7 100
16 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 40 95.8 0.0 0.0 53.3 99.3
17 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 20 73.6 0.0 0.0 66.7 100
18 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0 0.0 0.0 0.0 53.3 99.3
19 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 13.3 57.1 0.0 0.0 66.7 100
20 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0 0.0 0.0 0.0 40 95.8
V-B Finetuning
Fine-tuning LLMs is widely used to adapt models to specific downstream tasks [4, 3]. The model leverages knowledge acquired from pre-training on large unsupervised language learning tasks, allowing its weights to be adjusted using a smaller, task-specific dataset. This enables the model to capture novel patterns and improve task-specific reasoning without training from scratch. Fine-tuning LLMs offers several key advantages: (i) Reduced training time and computational costs by leveraging pre-trained foundational knowledge, (ii) Enhanced performance on downstream tasks, particularly beneficial when data is scarce, and (iii) Improved transfer learning across subtasks within a domain, increasing efficiency in model deployment.
Benchmarking LLMs: We evaluated the capabilities of LLMs in analog circuit design, using both open-source (CodeLlama-70B [39], DeepSeek-V2 [40]) and proprietary (GPT-3.5,4o-mini [41]) models. We fine-tuned all baseline LLMs using the corpus collected through the MASALA-CHAI framework. Open-source models were fine-tuned on a Nvidia A100 80GB GPU utilizing LoRA [42], while GPT models were fine-tuned using their respective fine-tuning playground [43].
[…]
Testing: We selected 20 design problems as shown in  Table II, frequently used as standalone circuits or subcomponents in larger designs, to form our evaluation benchmark. LLMs were prompted to generate the corresponding SPICE netlists.
Discussion: Table III summarizes the Pass@k performance of various LLMs for SPICE netlist generation. As anticipated, open-source models (DeepSeek and CodeLlama) underperform relative to the proprietary GPT models, both before and after fine-tuning. Nonetheless, fine-tuning on the collected dataset, (ID=1–5) provides a slight increase in the Pass@k for these open-source models, enabling coverage for additional design IDs (ID=6–8). Similarly, fine-tuned GPT-3.5 and GPT-4o achieve higher scores (ID=1–13) and non-zero performance for design IDs (ID=14–20). Notably, the fine-tuned GPT-4o (“Best Result”) successfully generates SPICE netlists for all design IDs, highlighting the efficacy of the MASALA-CHAI framework (also illustrated in Figure 1).
Figure 10 showcases representative examples of both successful and failed netlists generated by LLMs, illustrating critical design challenges and the effectiveness of our approach. Notably, in Figure 10: (1) demonstrates the improper use of an NMOS as a current source rather than forming the intended cascode configuration; (2) highlights a fundamental error where the gate of a diode-connected transistor, expected to connect to the drain, is incorrectly connected to the source; (3) the complexity of a two-stage design further amplifies the difficulties in understanding and replicating the intended topology. These examples show the limits of unoptimized LLMs. Through the finetuned GPT-4o model, we resolved the issues such as eliminating floating nets. For more analysis see Appendix A-C.
Figure 10: Pass/fail design cases generated by fine-tuned model.
V-C Integration into the AnalogCoder [14] Framework
AnalogCoder is a feedback-driven LLM framework for automated analog circuit design. Table IV reports the improvements in Pass@1 and Pass@5 on Tasks 12-20 by using our fine-tuned GPT-4o (Table III) with AnalogCoder instead of the base GPT-4o model (excluding tasks where AnalogCoder achieves 100% success). AnalogCoder’s Pass@1 rate on Task 20 improves from 20% to 60% when used with our fine-tuned GPT-4o model. The fine-tuned GPT-4o model alone only achieves a Pass@1 of 40%, suggesting that Masala-CHAI fine-tuning and AnalogCoder feedback together provide best performance.
TABLE IV: Results with AnalogCoder [14]. Many tasks show improvement and for others, results are the same (100%).
Task ID AnalogCoder AnalogCoder w/ FT GPT-4o
Pass@1 Pass@5 Pass@1 Pass@5
12 13.3 57.1 80 100
14 73.3 100 100 100
15 13.3 57.1 86.7 100
16 6.7 33.3 53.3 99.3
17 0.0 0.0 60 99.8
19 60 99.8 80 100
20 20 73.6 60 99.8
VI Conclusion
Masala-CHAI is an automated framework for creating large-scale SPICE netlist datasets from analog textbooks, resulting in the largest publicly available dataset of  7500 SPICE netlists with corresponding captions. Fine-tuning LLMs on Masala-CHAI demonstrates its use in automated SPICE netlist generation from natural language prompts. Our dataset and codebase (https://masala-chai-llm.github.io/) are fully open-sourced and easily extensible on additional textbooks and documentation.
Appendix A Appendix
A-A Evaluation
@@
We leverage graph-matching concepts to compare the structural similarity between SPICE netlists. Unlike traditional approaches, we parse the netlists to extract components and their connections. We identify components by types (e.g., Resistor, Capacitor, MOSFET) to ensure that our graph representation is invariant to component labels used during netlist generation. This yields a graph “G”, where nodes are circuit components and edges are interconnections, as illustrated Figure 11. We use Graph Edit Distance (GED) to compare representations of two SPICE netlists. GED ($G_{1}$,$G_{2}$) calculates the sum of the costs associated with transforming $G_{1}$ into $G_{2}$ through a series of edit operations on nodes and edges, such as insert, delete, and substitute. We normalize the GED to yield a final metric in the range 0 to 100. This normalization is achieved by calculating the maximum GED, obtained by summing the number of nodes and edges in both graphs and scaling accordingly. This metric allows for a consistent comparison across different circuits. Thus, the similarity score between the two netlists is:
@@
A-B Verifying against AMSNet [13] netlists
@@
We demonstrated the effectiveness of our approach by evaluating it using the AMSNet [13] dataset. Unlike AMSNet, which relies on an algorithmic method supplemented with some manual effort to generate SPICE netlists, our method introduces a fully automated flow (Figure 5). When comparing the SPICE netlists generated by our method to those produced by AMSNet, our approach achieved a 100% similarity score, as defined earlier. This validates the robustness of our flow and motivates us to extend the evaluation to additional circuit diagrams. Figure 11 illustrates an example of graphs representation of SPICE netlist generated by AMSNet (➊) and our Masala-CHAI framework (➋). The results demonstrate that the number of components and their neighboring elements is identical across both methods, with differences in component and net naming conventions. These findings strengthen Masala-CHAI accuracy and scalability in generating accurate netlists for various circuit diagrams.
A-C Role of high-quality dataset on Finetuning
Figure 12: Example showing responses from a fine-tuned GPT-4o model with AMSNet and MASALA-CHAI dataset, respectively.
As a part of our study, we fine-tuned the GPT-4o with the AMSNet [13] dataset as well and compared the quality of response with the MASALA-CHAI dataset. Figure 12 shows the responses for a simple “Common Source Amplifier” (ID=1) from the fine-tuned model using AMSNet and MASALA-CHAI dataset, respectively. AMSNet lacks a description of the SPICE netlist present in the dataset, which hinders its performance for supervised fine-tuning (SFT), making it not ready to use for LLM-based tasks that require high-quality description alongside the code for unlocking its full potential. MASALA-CHAI also offers such a pipeline to generate description alongside the SPICE netlist from a document, as shown Figure 5 (➑).